# Tilt your Head: Activating the Hidden Spatial-Invariance of Classifiers

### v1

@@ 5.2 Zero-Shot Affine Canonicalisation
Similar to STNs, we can undo affine transformations (i.e., canonicalize). We compare ITS to a STN (Jaderberg et al., 2015) with one and three transformation layers using inverse compositional layers (Lin & Lucey, 2017), a Diffeomorphic Transformer Network (DTN) (Detlefsen et al., 2018), and a Equivariant Transformer Network (ETN) (Tai et al., 2019). To study the impact of depth, we contrasted a $3$- and a $5$-layer ITS. We used a layer stack of rotation + scaling + shearing + (rotation + scaling). We refer to these as ITS 3 and ITS 5, respectively. Table 1 shows the quantitative results of our experiments. The associated qualitative results as reconstructions (after applying the predicted inverse transformation) can be found in Figure 8.
We found all baselines incapable of performing affine canonicalisations outside the training data distribution. This makes sense, as their localisation networks (regressing the transformation parameters) learn pixel-to-transformation mappings from training data. They are unable to generalise - particularly to rotations, which are mostly present during testing. The STN and the DTN estimate the composed transformation matrix in one forward pass. They need to extrapolate from nearly no rotations during training to $360^{\circ}$ rotations during testing. As the rotation manifold is isomorphic the unit circle ($\cos(\theta),\sin(\theta)$), this non-linearity needs to be learned implicitly. The ETN on the other hand uses canonical coordinates, which linearise the relationship between rotations.
5.3 Scaling up to ImageNet
@@
We found significant improvements on SI-Rotation as denoted in Table 2. To some extent this was expected as our confidence measure is highly sensitive in detecting canonical rotations (see Figure 3). Surprisingly, however, our method even outperformed augmented fine-tuning and explicit training on the foreground object (Chefer et al., 2022). On SI-Size ITS maintains the performance of the baseline. We hypothesise that location and size are properties that have high variance in the vanilla ImageNet already. This puts the test performance on a high level without requiring any additional mechanisms. Equipping the backbone with the ability to translate the input query with two degrees of freedom causes it to focus on the wrong object. This behaviour is similar to attention but with a fixed centre bias of the backbone and a spatial translation of the signal instead. However, ITS does not get any feedback (gradient) if the object of focus is correct. This leads to slightly degenerated performance on SI-Location (see Section B.3).
@@ 6 Conclusion
Inverse Transformation Search (ITS) equips any vanilla image classifier with pseudo-invariance to various spatial transformations during inference. This increases its robustness to spatially perturbed inputs without requiring extra training data. We activate this potential, which lies hidden in any classifier, without requiring any inductive biases or augmented training. This is achieved by an incremental search for the inverse perturbation. At the canonical form, the features of the transformed input match the features known from the training data. Here the model is most confident to classify this sample properly. To mitigate the issue of spatial collapses during repetitive transformations, we propose a group-induced confidence measure. We successfully showed the superiority of our methods for zero-shot image classification tasks.
Future Work
One issue is the rigidness of existing image datasets. Heavily perturbed images can have multiple proper hypotheses. See the example in Figure 6, where both labels ($5$ or $9$) seem reasonable. As ITS uses the representation space of the trained classifier, its performance depends on the quality of these representations. The model might detect features in transformed signals, which lead to false hypotheses. Even worse, these hypotheses might be reinforced down the search tree. Easing these instabilities should be the main objective of future investigations.
@@ Appendix A Implementation Details
For educational purposes, we used $M=1$ (i.e., a deterministic search), which ensures consistently improving scores along the paths in Figures 5 and 6. We used a group cardinality of $n=17$ for all our experiments due to algorithmic reasons. We divide the parameter space $[-\theta,\theta]$ by $n$ into an equally distant parameter set. The identity $\theta=0$ lies in the middle of this domain and is therefore only included for odd $n$. For instance, $n=7$ gives $3$ clockwise rotations, $3$ counter-clockwise rotations and the identity (no rotation).
A.1 Model Architectures
@@
This is used for experiments on MNIST and Fashion-MNIST. This model comprises three 2D-convolutional layers with $5\times 5$ kernels and $8$, $16$, and $16$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2023) after each convolution layer. 2D Max-Pooling with $2\times 2$ kernels is added after the second and the third convolutional layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
@@
This is used for experiments on GTSRB. This model comprises three 2D-convolutional layers with kernel sizes of $5$, $3$ and $1$ and $100$, $150$, and $250$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2023), 2D Max-Pooling with $2\times 2$ kernels, Batch Normalisations (Ioffe & Szegedy, 2015) and Dropout after each convolution layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
Appendix B Further Experiments
B.1 Runtime
@@
All experiments are performed on an Nvidia A40 GPU (48GB) node with 1 TB RAM, 2x 24core AMD EPYC 74F3 CPU @ 3.20GHz, and a local SSD (NVMe). The software specifications of our implementations can be found in our open-sourced code. Figure 9 shows how the runtime behaves under different hyper-parameter settings compared to the pure backbone. The most significant factor is the depth of the search tree. In Section 5.2 we found that the performance does not notably increase with more hypotheses. The number of Monte-Carlo dropout steps, $M$, stabilises the predictions by reducing temporal confidence anomalies. Although, we used $M=50$ for our experiments, the impact on the performance is minor. Another factor is the number of hypotheses tested. These hypotheses are processed in parallel but due to the unique class condition (see Section 4), the runtime scales with $k$. Empirically, we found that $k=2$ produces the best results, which we used in all of our experiments. Furthermore, we want to emphasise again that our inference strategy is targeted for applications where it is worth trading runtime against robustness, as in (Ghamizi et al., 2022).
B.2 Change of Mind
@@
Appendix C Theoretical Specifications
C.1 Transformation Decomposition
@@
C.2 Convolution over Finite Orbits
@@
Lemma C.1 (Convolution over Finite Orbits).
@@
C.3 Markov Interpretation
@@
More specifically, the search describes a Markov process. The Markov property of a stochastic process guarantees that the probability of a future state depends only on the current state independently of its past behaviour. This implies that the branching candidate in each level only depends on the previous candidate. This is true, as every roll-out depends only on the prior candidate in the first place. These one-step dynamics simplify the system. In Section D.4 we provide a sketch of the proof.
Appendix D Proofs
D.1 Proof: Lemma C.1
@@ Appendix D Proofs
Let $E_{\mathbf{\theta}}(\mathbf{x})$ be the energy of a $\theta$-parameterised energy model given a point $\mathbf{x}$ as input as defined in Equation 4. Let $\bar{E}_{\mathbf{\theta}}(\mathbf{x})$ be the Monte-Carlo estimate of the energy using Equation 5. As shown in Lemma C.1 we can convolve this energy function on the orbit with a kernel, $K$. To remove visual clutter, we substitute
@@
Given a Matrix Lie Sub-Group, $H$ and a finite sub-group, $G\leq H$. We can approximate $H$ by increasing the cardinality of $G$, such that in the limit $G=H$. Let $\langle g\rangle$ be the generator of $G$, which forms an atomic element of $G$. Then, every $g\in G$ can be expressed by $g=\hat{g}^{k}$ for some $k$. Using composition, a step along the orbit starting at $g$ can be expressed by $(\hat{g}\circ g)\mathbf{x}$. Similarly, a step in the opposite direction can be performed by $(\hat{g}^{-1}\circ g)\mathbf{x}$. We substitute $g\mathbf{x}+h$ by $(\hat{g}\circ g)\mathbf{x}$ in Section D.2.
Let $\mathbf{T}_{g}$ be the matrix representation of $g$. The step length along the orbit is defined by $\|\mathbf{T}_{\hat{g}}-\mathbf{T}_{\hat{g}^{2}}\|$. These two observations lead to the following reformulation of Section D.2,
@@
D.4 Proof: Markov Process
@@
Definition D.1 (Stochastic Process (Borodin, 2017)).
@@
Definition D.2 (Markov Process (Borodin, 2017)).

### v2

@@ 5.3 Scaling up to ImageNet
Figure 9: SI-Rotation sample, showing a tilted image of a cropped cab (ground-truth label) pasted on a mountain-view background. We applied ITS with rotation $[-\pi,\pi]$ (first row) and isotropic-scaling $[-\frac{1}{2},\frac{1}{2}]$ (second and third row) with two hypotheses ($H_{0}$ and $H_{1}$).
@@
We found significant improvements on SI-Rotation as denoted in Table 2. To some extent this was expected as our confidence measure is highly sensitive in detecting canonical rotations (see Figure 3). Surprisingly, however, our method even outperformed augmented fine-tuning and explicit training on the foreground object (Chefer et al., 2022). An example is depicted in Figure 9. This shows the brittleness of ImageNet-trained models to spatial transformed inputs (Engstrom et al., 2019) (especially to rotations), as the label is changing frequently along the orbits. An exception is $H_{1}$ (third row), where the correct label is preserved over almost the entire orbit of isotropic-scalings. This property is studied in SI-Size.
On SI-Size ITS maintains the performance of the baseline. We hypothesise that location and size are properties that have high variance in the vanilla ImageNet already. This puts the test performance on a high level without requiring any additional mechanisms. Equipping the backbone with the ability to translate the input query with two degrees of freedom causes it to focus on the wrong object. This behaviour is similar to attention but with a fixed centre bias of the backbone and a spatial translation of the signal instead. However, ITS does not get any feedback (gradient) if the object of focus is correct. This leads to slightly degenerated performance on SI-Location (see Section C.4).
Furthermore, we tested two variants to obtain acc@5: (i) We computed $H_{0}$’s categorical distribution and picked the top-5 predictions. (ii) We used the top-1 predictions of each of the five hypotheses ($H_{0}\ldots H_{4}$). A higher score for the former would indicate that on average $H_{0}$ followed the correct class hypothesis. Whereas, when the latter is higher, one of the alternative hypotheses ($H_{1}\ldots H_{4}$) contained the true class more frequently than $H_{0}$. Unfortunately, our experiments do not show a consistent pattern of which strategy yields higher accuracies. We leave further investigations to future work.
@@ 6 Conclusion
Inverse Transformation Search (ITS) equips any vanilla image classifier with pseudo-invariance to various spatial transformations during inference. This increases its robustness to spatially perturbed inputs without requiring extra training data. We activate this potential, which lies hidden in any classifier, without requiring any inductive biases or augmented training. This is achieved by an incremental search for the inverse perturbation. At the canonical form, the features of the transformed input match the features known from the training data. At these points, the model is most confident to classify this sample properly. We successfully showed the superiority of our methods for zero-shot image classification tasks.
6.1 Limitations
Mitigating the following three limitations should be the focus of future works: (1) False Hypotheses: Heavily perturbed images can have multiple proper hypotheses. See the example in Figure 6, where both labels ($5$ or $9$) seem reasonable. As ITS uses the representation space of the trained classifier, its performance depends on the quality of these representations. The model might detect features in transformed signals, which lead to false hypotheses. Even worse, these hypotheses might be reinforced down the search tree. See Figure 12 for an example of such behaviour. (2) Abelian Assumption: In Theorem 4.1, we showed that only a subsequence of group actions need to be Abelian. Although, this alleviates the impact of the assumption, the issue is not solved. (3) Runtime Cost: Due to the complexity of the search process (see Section 4) the walltime increases significantly during inference (see Section C.1). In many setups, like video understanding (Bexten et al., 2021), this renders real-time applications intractable.
6.2 Future Work
Given the energy field of the target model, any post-hoc optimisation method can be applied to maximise the confidence of a test sample over a transformation group. This search can be done over the entire group space, not only over subgroups (as used here) tackling the issue with Abelian assumption. This opens up a wide range of algorithms, including for example simulated annealing, which is used in a wide range of domains (Landau & Binder, 2005; Schmidt et al., 2024). As such algorithms are applicable to Quantum-Computers, they hold the potential to eliminate the runtime issue.
@@ 7 Impact Statement
8 Acknowledgement
Special thanks to all our lab members for the extremely valuable feedback and support during the process. Additionally, we thank the anonymous reviewers for their insightful comments that helped improve this work.
@@ Appendix A Implementation Details
If not otherwise stated in the experiment descriptions, we used a group cardinality of $n=17$. The choice of odd numbers is due to our algorithmic implementation. We divide the parameter space into $n$ equally distant chunks. An odd $n$ ensures that the identity is included.
A.1 Search Algorithm
Supplementary to the descriptions and illustrations in Figure 4 and Figure 5, we provide the pseudocode of our proposed algorithm in Algorithm 1. The complexity of the search increases if more than one hypothesis is maintained. We refer the reader to our publicly available source code for more details.444www.github.com/johSchm/ITS
Algorithm 1 Pseudocode for ITS with one hypothesis $H$.
  Input: $\mathbf{x}\sim G\mathcal{D}_{\text{test}}$
  $\rho_{\text{primal}}=0$ and $h=\{e\}$
  for $G$ in $\{G\}$ do
     $\rho,g=\arg\max_{g\in G}[\rho_{G}(\mathbf{x},g\circ h)]$
     if $\rho>\rho_{\text{primal}}$ then
        $\rho_{\text{primal}}\leftarrow\rho$
        $H\leftarrow p_{\theta}(y\mid g\circ h\mathbf{x})$
        $h\leftarrow g\circ h$
     end if
  end for
  Output: $h$
A.2 Model Architectures
@@
This is used for experiments on MNIST and Fashion-MNIST. This model comprises three 2D-convolutional layers with $5\times 5$ kernels and $8$, $16$, and $16$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2016) after each convolution layer. 2D Max-Pooling with $2\times 2$ kernels is added after the second and the third convolutional layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
@@
This is used for experiments on GTSRB. This model comprises three 2D-convolutional layers with kernel sizes of $5$, $3$ and $1$ and $100$, $150$, and $250$ channels. We add GeLU non-linearities (Hendrycks & Gimpel, 2016), 2D Max-Pooling with $2\times 2$ kernels, Batch Normalisations (Ioffe & Szegedy, 2015) and Dropout after each convolution layer. The final stack of feature maps is flattened into a vector and further compressed to the number of classes by two linear layers. After the first linear layer, we also add a GeLU non-linearity and a dropout-layer with $50\%$ of dropout. From here we read out the logit scores.
Appendix B Further Related Works
B.1 Test-time Domain Adaptation
In our work, we model a canonicalisation function by a search process. Canonicalisation (Kaba et al., 2023; Mondal et al., 2023) is build on first principles and supported by the mathematical guarantees of group theory. This is highly distinctive to learnable test-time adaptions (TTAs) (Wang et al., 2021, 2022; Wu et al., 2024). We do not need to modify the parameters of the neural network, we can even treat the model as a black box only having access to the logit-scores. Source-free TTAs often suffer from error accumulations and catastrophic forgetting (Wang et al., 2022). For our method, the latter is avoided a priori, as at no point the neural network and its weights are modified. Similar to (Wang et al., 2022), we use weight averages (approx. Bayesian) to mitigate error accumulations.
Appendix C Further Experiments
Remark regarding Image-Backgrounds
Spatial transformations can introduce background artefacts for datasets without mono-colour backgrounds, like ImageNet. There exist only solutions for a few quotient groups (like rotations (Tuggener et al., 2023)) not for the entire affine group (as considered in this work). Fortunately, this is only an issue during training as the model might exploit these artefacts, which does not apply for any conducted experiment. Either all models are trained on the vanilla dataset (as in Section 5.1 and Section 5.2) or are trained with artefacts (baselines in Section 5.3) but tested on an artefact-free dataset (here SiScore).
C.1 Runtime
@@
All experiments are performed on an Nvidia A40 GPU (48GB) node with 1 TB RAM, 2x 24core AMD EPYC 74F3 CPU @ 3.20GHz, and a local SSD (NVMe). The software specifications of our implementations can be found in our open-sourced code. Figure 10 shows how the runtime behaves under different hyperparameter settings compared to the pure backbone. The most significant factor is the depth of the search tree. In Section 5.2 we found that the performance does not notably increase with more hypotheses. The number of Monte-Carlo (MC) dropout steps, stabilises the predictions by reducing temporal confidence anomalies. Although, we used $50$ MC samples for our experiments, the impact on the performance is minor. Another factor is the number of hypotheses tested. These hypotheses are processed in parallel but due to the unique class condition (see Section 4), the runtime scales with the number of hypotheses. Empirically, we found that three to five produces the best results.
Figure 11: Average acc@1 improvement on the randomly rotated and scaled MNIST test set compared to the baseline (without ITS) performance.
C.2 Hyperparameter Tuning for Accuracy
We continue the previous investigation by studying the influence of these hyperparameters on the accuracy. Increasing the number of MC samples stabilises predictions but also reduces exploration. When increasing the MC samples from $1$ to $10$ on affine-transformed Fashion-MNIST the acc@1 increases $9.46\pm 1.14$ points (mean and standard deviation over $10$ runs). Increasing the MC samples from $10$ to $50$ leads to an acc@1 increase of only $1.38\pm 0.47$.
In Figure 11, we illustrated the performance improvements with respect to the number of samples per layer and the number of hypotheses. Interestingly, the performance does not increase (or stagnate) when increasing either value. It seems that the middle ground around $|H|=3$ and $|G|=16$ results in maximum improvements compared to the vanilla model (without ITS) prediction. Due to the unique-label constraint of each hypothesis (see Section 4) many hypotheses with initial low confidence are pursued. This might lead to catastrophic reinforcement of many false hypotheses, obtaining higher confidences down the line.
C.3 Change of Mind
@@
model acc@1 acc@5
SI-Rotation EfficientNet 25.1 43.0
EfficientNet + ITS 34.4 (+9.3) 47.9 (+4.9)
ConvNeXt 42.4 65.3
ConvNeXt + ITS 54.7 (+12.3) 70.0 (+4.7)
SI-Size EfficientNet 35.7 57.3
EfficientNet + ITS 36.9 (+1.2) 50.5 (-6.8)
ConvNeXt 53.2 75.8
ConvNeXt + ITS 56.4 (+3.2) 68.3 (-7.5)
SI-Location EfficientNet 15.7 30.3
EfficientNet + ITS 11.4 (-4.3) 17.9 (-12.4)
ConvNeXt 32.1 58.6
ConvNeXt + ITS 31.6 (-0.5) 51.5 (-7.1)
Table 4: Continuation of the SiScore experiment with an EfficientNet-B0 (Tan & Le, 2019) and a ConvNeXt-tiny (Liu et al., 2022) backbone.
Furthermore, we tested other popular backbone architectures, covering a EfficientNet (Tan & Le, 2019) and a ConvNeXt (Liu et al., 2022). Both models were pretrained on ImageNet-1K. The results are shown in Table 4.
Appendix D Theoretical Specifications
D.1 Transformation Decomposition
@@
D.2 Convolution over Finite Orbits
@@
Lemma D.1 (Convolution over Finite Orbits).
@@
The proof can be found in Section E.1.
D.3 Markov Interpretation
@@
More specifically, the search describes a Markov process. The Markov property of a stochastic process guarantees that the probability of a future state depends only on the current state independently of its past behaviour. This implies that the branching candidate in each level only depends on the previous candidate. This is true, as every roll-out depends only on the prior candidate in the first place. These one-step dynamics simplify the system. In Section E.4 we provide a sketch of the proof.
Appendix E Proofs
E.1 Proof: Lemma D.1
@@ Appendix E Proofs
Let $E_{\mathbf{\theta}}(\mathbf{x})$ be the energy of a $\theta$-parameterised energy model given a point $\mathbf{x}$ as input as defined in Equation 4. Let $\bar{E}_{\mathbf{\theta}}(\mathbf{x})$ be the Monte-Carlo estimate of the energy using Equation 5. As shown in Lemma D.1 we can convolve this energy function on the orbit with a kernel, $K$. To remove visual clutter, we substitute
@@
Given a Matrix Lie Sub-Group, $H$ and a finite sub-group, $G\leq H$. We can approximate $H$ by increasing the cardinality of $G$, such that in the limit $G=H$. Let $\langle g\rangle$ be the generator of $G$, which forms an atomic element of $G$. Then, every $g\in G$ can be expressed by $g=\hat{g}^{k}$ for some $k$. Using composition, a step along the orbit starting at $g$ can be expressed by $(\hat{g}\circ g)\mathbf{x}$. Similarly, a step in the opposite direction can be performed by $(\hat{g}^{-1}\circ g)\mathbf{x}$. We substitute $g\mathbf{x}+h$ by $(\hat{g}\circ g)\mathbf{x}$ in Section E.2.
Let $\mathbf{T}_{g}$ be the matrix representation of $g$. The step length along the orbit is defined by $\|\mathbf{T}_{\hat{g}}-\mathbf{T}_{\hat{g}^{2}}\|$. These two observations lead to the following reformulation of Section E.2,
@@
E.4 Proof: Markov Process
@@
Definition E.1 (Stochastic Process (Borodin, 2017)).
@@
Definition E.2 (Markov Process (Borodin, 2017)).